Given the demand function :
P = 500 - 4Q^2
Calculate arc elasticity averaged along the arc joining Q = 8 and Q = 10
1
Expert's answer
2018-04-13T15:50:14-0400
E = (Q_2 - Q_1)*(P_1+P_2) / ((Q_2 + Q_1)*(P_2-P_1))
Q_1 = 8
Q_2 = 10
P_1 = 500 - 4*8^2 = 244
P_2 = 500 - 4*10^2 = 100
Then
E = (10 - 8)*(244+100) / ((10 + 8)*(100-244)) = - 2*344/(18*144) = -0.27
Demand is relatively inelastic